If a firm has a price of earnings of 9, a market to book ratio of 2.5 and an EPS of 3.50 what is the book to value per share? Please show work
Answer #1

In my understanding, the question asks to calculate Book Value per share (BVPS) .

Given that Price of Earnings, that is, P/E Ratio is 9. Also that, EPS (Earnings per share) = 3.50

P/E Ratio = MPS / EPS = Market Price per share / Earnings per share

=> 9 = MPS / 3.50

=>MPS = 9*3.50 = 31.50

Now, it's given that Market to Book Ratio is 2.5

Market to Book Ratio = Market Capitalisation of Company / Book Value of Company = Market Price per Share / Book Value Per share

Market to Book Ratio = MPS / BVPS

=> 2.50 = 31.50 / BVPS

=> BVPS = 31.50 / 2.50 = 12.6

So, Book Value per Share (BVPS) is 12.60

Note:

  • We have made all the calculations in per-share numbers. If you face any question where the Book Value of company is asked or the Market capitalisation of the company is asked, you can use the alternate formula ( Market to Book Ratio = Market Capitalisation of Company / Book Value of Company) as mentioned above.
  • Price to Earnings or P/E Ratio states the no. of times of EPS that an investor is ready to pay to buy the stock to earn an amount equivalent to the EPS. Usually, higher EPS suggests more confidence of the investor in the company's future growth.